OpenStudy (anonymous):

how would you simplify ²√192¹

OpenStudy (anonymous):

\[2\sqrt{192^1}\]like this?

OpenStudy (anonymous):

or\[\sqrt{192^1}\]

OpenStudy (anonymous):

no the 2 is an exponent

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[192^{(\frac{ 1 }{ 2 }\times2)\times1}\]

OpenStudy (anonymous):

\[\sqrt{}=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

not understanding

OpenStudy (anonymous):

which part u don't understand?

OpenStudy (anonymous):

I divided 2 into 192 and got 96...but 2 can go into 96 and I got 48....2 can go into 48...and I got 24....2 can go into 24....12...6.....3...so why wouldn't it be 1/3

OpenStudy (anonymous):

First,convert the square root into exponent

OpenStudy (anonymous):

\[\sqrt{}=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

so 1/2 is the exponent

OpenStudy (anonymous):

\[192^{(\frac{ 1 }{ 2 }\times2)\times1}\]